# Adaptive and Efficient Learning with Blockwise Missing and Semi-Supervised Data

### v1

@@ 5 Heart Disease Risk Modeling with MIMIC-III
To illustrate the utility of DEFUSE, we apply it to MIMIC-III (Medical Information Mart for Intensive Care III; Johnson et al., (2016)), a large database comprising de-identified electronic health records (EMRs) of over forty thousand patients associated with their stay in critical care units of the Beth Israel Deaconess Medical Center between 2001 and 2012. The database includes various types of information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality in and out of the hospital. As EMRs include large and diverse patient populations and cover comprehensive phenotypes, they have been widely used by researchers to identify underlying risk factors for certain disease conditions. Our goal is to make statistical inference on the risk prediction model for heart diseases (HD) against some potential risk factors related to the health condition of a subject.
@@
Our response of interest $Y\in\{0,1\}$ is the HD status identified from the manual chart reviewing labels obtained in a previous study on MIMIC-III data (Gehrmann et al.,, 2018); see their paper for the detailed definition of the phenotype HD. We greatly thank the authors for making their gold standard labels accessible. For $\boldsymbol{X}$, we include age, encounter of the International Classification of Diseases (ICD) codes of low high-density lipoproteins (HDL), ICD encounter of type II diabetes, and several HD-related laboratory (lab) variables selected using KESER (Hong et al.,, 2021), an online tool for extracting clinical concepts and features relevant to a given phenotype. We exclude lab variables too closely linked with HDL or diabetes such as A1C. For the convenience of downstream risk modeling, we also derive a lab risk score (LRS) through weighting and combining the selected lab variables with the main ICD code of HD as a surrogate outcome on the large $\mathcal{U}\mathcal{C}$ sample. Since the LRS is derived with large sample, it has an negligible estimation uncertainty and can be viewed as given. See the list of selected lab variables and details about the LRS in Appendix.
@@
Noticing that each lab variable is not observed on a fraction of patients, we separately consider two settings of missingness: (i) a single BM task with one BM covariate taken as the derived LRS; (ii) a multiple BM task with two BM lab variables thyroid-stimulating hormone (TSH) and hypochromia (HPO), with joint missingness on $\mathcal{L}\mathcal{M}_{1}$, TSH-only missingness on $\mathcal{L}\mathcal{M}_{2}$, HPO-only missing on $\mathcal{L}\mathcal{M}_{3}$. On task (i), there are $N_{\mathcal{L}\mathcal{C}}=467$ labeled subject with complete observation of $\boldsymbol{X}$, $N_{\mathcal{L}\mathcal{M}}=678$ labeled subjects with BM covariates, and $N_{\mathcal{U}\mathcal{C}}=7037$ unlabeled subjects with complete $\boldsymbol{X}$. Similar to the simulation, we compare our approaches to four applicable benchmarks to this case, including $\mathcal{L}\mathcal{C}$-only, SSL, Semieff, and HTLGMM. We do not include MICE since it tends to result in excessive bias, as seen from the simulation results. For (ii), there are $N_{\mathcal{L}\mathcal{C}}=486$ labeled subject with complete observation of $\boldsymbol{X}$, $N_{\mathcal{L}\mathcal{M}_{1}}=231$, $N_{\mathcal{L}\mathcal{M}_{2}}=175$, and $N_{\mathcal{L}\mathcal{M}_{3}}=153$ for three types of missing, and $N_{\mathcal{U}\mathcal{C}}=7014$. For the implementation of DEFUSE, we follow the same strategy as our simulation with the conditional models constructed through kernel machine and the data-adaptive calibration functions $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ and $h(\boldsymbol{x},y;\boldsymbol{\zeta})=1+\zeta_{1}$.
Table 6 provides a summary of the estimated coefficients on the single BM task (i), as well as the bootstrap standard errors and $p$-values. Table 7 reports the relative efficiency (RE) of the data-fusing estimators to the standard $\mathcal{L}\mathcal{C}$-only estimator. DEFUSE achieves the highest RE on all the coefficients among the included methods. As an example, for the estimation of the HDL coefficient, our method reduces more than $50\%$ variance compared to $\mathcal{L}\mathcal{C}$-only and its RE to HTLGMM is around $1.4$.
@@
Through the output of DEFUSE, one could conclude that given the age and lab measures on other risk factors, a higher risk of HD is associated with having low HDL at the level $0.1$ as well as type II diabetes at the level $0.05$. Notably, with the methods $\mathcal{L}\mathcal{C}$-only and SSL, both the effects of HDL and diabetes are shown to be insignificant even at the $0.1$ level. Also, diabetes turns out to be not significant when using HTLGMM. This suggests that among all the methods under comparison, DEFUSE is the only one powerful enough to discover both signals. Interestingly, it is well-documented in existing biomedical literature on the associations between heart disease and diabetes (Haffner,, 2000; Peters et al.,, 2014, e.g.) as well as low HDL (Després et al.,, 2000; Rader and Hovingh,, 2014, e.g.), which supports the findings of DEFUSE. Table 8 summarizes the RE to the $\mathcal{L}\mathcal{C}$-only estimator in the multiple BM scenario introduced above. DEFUSE attains more than $25\%$ efficiency improvement over $\mathcal{L}\mathcal{C}$-only on all coefficients and more than $50\%$ variance reduction on that of age, which is again more significant than SSL. One can find the summary table of the estimates and $p$-values in this scenario in Appendix.
@@
$\mathcal{L}\mathcal{C}$-only $0.41_{0.120}$ $<$0.001 $0.35_{0.219}$ 0.120 $0.15_{0.215}$ 0.48 $0.25_{0.113}$ 0.028
DEFUSE1 $0.46_{0.0823}$ $<$0.001 $0.25_{0.148}$ 0.086 $0.34_{0.146}$ 0.020 $0.22_{0.112}$ 0.048
DEFUSE $0.45_{0.0821}$ $<$0.001 $0.25_{0.147}$ 0.087 $0.33_{0.145}$ 0.022 $0.22_{0.111}$ 0.051
SSL $0.41_{0.140}$ 0.003 $0.36_{0.240}$ 0.129 $0.15_{0.238}$ 0.532 $0.25_{0.113}$ 0.027
HTLGMM $0.47_{0.102}$ $<$0.001 $0.37_{0.192}$ 0.053 $0.20_{0.193}$ 0.295 $0.25_{0.117}$ 0.032
Table 6: Fitted coefficient (coe), empirical standard error (se), and $p$-value for the coefficient of each covariate in the single BM scenario of our real-world study.
Method Age HDL Diabetes Lab Score
DEFUSE1 2.11 2.19 2.18 1.00
DEFUSE 2.13 2.21 2.19 1.10
SSL 1.00 1.01 1.00 1.01
HTLGMM 1.89 1.60 1.47 1.00
Table 7: Relative efficiency (RE) to the standard $\mathcal{L}\mathcal{C}$-only estimator on each coefficient in the single BM scenario of our real-world study.

### v2

@@ 5 Heart Disease Risk Modeling with MIMIC-III
5.1 Extension on Auxiliary Covariates
We begin with introducing a natural extension of our framework to include auxiliary features, which will be used in our real-world study. Consider the scenario with covariates $\{\boldsymbol{X},\boldsymbol{W}\}$ consisting of the primary risk factors $\boldsymbol{X}$ and auxiliary or surrogate features $\boldsymbol{W}$. Instead of the whole $\{\boldsymbol{X},\boldsymbol{W}\}$, the focus is still on $Y\sim\boldsymbol{X}$ defined by the population equation (2). The missing structure on $(Y,\boldsymbol{X})$ remains the same as described in Section 1.2, i.e., one or multiple subsets of covariates $\boldsymbol{X}_{\mathcal{P}_{m}}$ partially missing on the labeled sample and $Y$ being unlabeled on $\mathcal{U}\mathcal{C}$ with complete $\boldsymbol{X}$. Meanwhile, suppose $\boldsymbol{W}$ to be complete on all subjects.
Such auxiliary features $\boldsymbol{W}$ are commonly used in practice. For instance, in EMR linked biobank study, the interest lies in the relationship between certain genetic variants $\boldsymbol{X}$ and disease status $Y$. $\boldsymbol{W}$ is taken as EMR proxies (e.g., relevant diagnostic codes) for $Y$ and acts as auxiliary features not included in the genetic risk model. In clinical or epidemiological studies, $\boldsymbol{X}$ is the treatment or key risk factors, $Y$ is a long-term outcome, and $\boldsymbol{W}$ can be some early endpoints or surrogate outcomes (e.g., tumor response rates) occurring and observed post the baseline with only $\boldsymbol{X}$. In this case, one is interested in $Y\sim\boldsymbol{X}$ to make clinical decision at the baseline without $\boldsymbol{W}$. In both examples, the auxiliary $\boldsymbol{W}$ could still be informative to $Y$ and $X$ and, thus, included to boost the statistical efficiency.
Our framework can be naturally extended to accommodate this setting. For the construction described in Section 2 and summarized in Algorithm 1, the only change to make is to allow the inclusion of $\boldsymbol{W}$ in the nuisance imputation models ${\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},\boldsymbol{W},Y;\boldsymbol{\eta})$ and $\nu(Y|\boldsymbol{X},\boldsymbol{W};{\boldsymbol{\theta}})$, as well as in the calibration functions $w(\boldsymbol{X},\boldsymbol{W},Y;{\boldsymbol{\delta}})$ and $h(\boldsymbol{X},\boldsymbol{W},Y;{\boldsymbol{\zeta}})$. In this case, $\boldsymbol{W}$ could often be of high-dimensionality as considered in Zhou et al., (2024) as well as our real example. The theoretical results established in Section 3 can also be applied to this setting. Despite the association of $\boldsymbol{W}$ with both $\boldsymbol{X}$ and $Y$, including $\boldsymbol{W}$ in the nuisance imputation models will not create bias thanks to our constructions like (7) enabling automatic bias correction with respect to the nuisance estimators. It is also important to note that our data-adaptive control function approach can be particularly useful in the presence of erroneous nuisance estimators with high-dimensional $\boldsymbol{W}$, which will be illustrated though the real example in this section.
5.2 Setup
@@
To illustrate the utility of DEFUSE, we apply it to MIMIC-III (Medical Information Mart for Intensive Care III; Johnson et al., (2016)), a large database comprising de-identified electronic health records (EMRs) of over forty thousand patients associated with their stay in critical care units of the Beth Israel Deaconess Medical Center between 2001 and 2012. The database includes various types of information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality in and out of the hospital. As EMRs include large and diverse patient populations and cover comprehensive phenotypes, they have been widely used by researchers to identify underlying risk factors for certain disease conditions. Our goal is to make statistical inferences on the risk prediction model for heart diseases (HD) against some potential risk factors related to the health condition of a subject.
@@
Our response of interest $Y\in\{0,1\}$ is the HD status identified from the manual chart reviewing labels obtained in a previous study on MIMIC-III data (Gehrmann et al.,, 2018); see their paper for the detailed definition of the phenotype HD. We greatly thank the authors for making their gold-standard labels accessible. For $\boldsymbol{X}$, we include age, encounter of the International Classification of Diseases (ICD) codes of low high-density lipoproteins (HDL), ICD encounter of type II diabetes, and several HD-related laboratory (lab) variables selected using KESER (Hong et al.,, 2021), an online tool for extracting clinical concepts and features relevant to a given phenotype. For the convenience of downstream risk modeling, we derive a lab risk score (LRS) through weighting and combining the selected lab variables with the main ICD code of HD as a surrogate outcome on the large $\mathcal{U}\mathcal{C}$ sample. Since the LRS is derived with a large sample, it has a negligible estimation uncertainty and can be viewed as given. See the list of selected lab variables and details about the LRS in Appendix.
@@
Noticing that each lab variable is not observed on a particular fraction of patients, we separately consider two settings of missingness: (i) a single BM task with one BM covariate taken as the derived LRS; (ii) a multiple BM task with two BM lab variables thyroid-stimulating hormone (TSH) and hypochromia (HPO), with joint missingness on $\mathcal{L}\mathcal{M}_{1}$, TSH-only missingness on $\mathcal{L}\mathcal{M}_{2}$, HPO-only missing on $\mathcal{L}\mathcal{M}_{3}$. On task (i), there are $N_{\mathcal{L}\mathcal{C}}=467$ labeled subject with complete observation of $\boldsymbol{X}$, $N_{\mathcal{L}\mathcal{M}}=678$ labeled subjects with BM covariates, and $N_{\mathcal{U}\mathcal{C}}=7037$ unlabeled subjects with complete $\boldsymbol{X}$. Similar to the simulation, we compare our approaches to four applicable benchmarks to this case, including $\mathcal{L}\mathcal{C}$-only, SSL, Semieff, and HTLGMM. We do not include MICE since it tends to result in excessive bias, as seen from the simulation results. For (ii), there are $N_{\mathcal{L}\mathcal{C}}=486$ labeled subject with complete observation of $\boldsymbol{X}$, $N_{\mathcal{L}\mathcal{M}_{1}}=231$, $N_{\mathcal{L}\mathcal{M}_{2}}=175$, and $N_{\mathcal{L}\mathcal{M}_{3}}=153$ for three types of missing, and $N_{\mathcal{U}\mathcal{C}}=7014$.
To assist data-fusion, we include auxiliary features $\boldsymbol{W}$ as introduced in Section 5.1. With all codified features including the ICD and procedure codes in EMR, we apply a simple pre-screening procedure with the ICD count of HD as a surrogate outcome against around all the EMR features on the whole sample, to select $140$ features and include them in $\boldsymbol{W}$ for the downstream risk model estimation. For the implementation of DEFUSE, we follow the same strategy as our simulation with the conditional models constructed separately through kernel machine (KM) and Lasso with their hyper-parameters tuned by cross-validation. Similar to our simulation design, the calibration functions are set as $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ and $h(\boldsymbol{x},y;\boldsymbol{\zeta})=1+\zeta_{1}$ and we also include DEFUSE1 with $w(\boldsymbol{x},y;\boldsymbol{\delta})=1$ for comparison with existing calibration strategies (Miao et al.,, 2023, e.g.). To ensure fair comparison, we use the same machine learning estimators for missing variable imputation in all the benchmark methods if needed.
5.3 Results
Table 7 reports the relative efficiency (RE) of the data-fusing estimators to $\mathcal{L}\mathcal{C}$-only in the single BM task (i). DEFUSE achieves the highest REs on all coefficients. For example, in estimating the HDL coefficient, our method reduces more than $50\%$ variance compared to $\mathcal{L}\mathcal{C}$-only and its RE to HTLGMM is around $1.41$. Interestingly, under both the KM and Lasso implementation, DEFUSE1 with the calibration basis $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ attains smaller estimation variance on the partially missing predictor LRS compared to SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$ while they all have close performance on the other predictors. In specific, with Lasso imputation, DEFUSE1 with $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ has $10\%$ improved efficiency over $\mathcal{L}\mathcal{C}$-only while either SemiEff or DEFUSE1 with $\xi(\boldsymbol{x},y)=1$ has no improvement over $\mathcal{L}\mathcal{C}$-only. Similarly, when using KM, DEFUSE1 with $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ attains a $9\%$ higher RE than SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$. This demonstrates not only the importance of data-adaptive calibration on the erroneous machine learning estimators, but also the effectiveness of our newly proposed control function strategy in Section 2.2 that is more general and sophisticated than the simple one with $\xi(\boldsymbol{x},y)=1$. In addition, on LRS, DEFUSE with the SSL step shows a further improvement over DEFUSE1 with 6-7% higher efficiency while such an improvement is not significant on the other coefficients.
@@ 5.3 Results
Table 8 provides a summary of the estimated coefficients as well as the bootstrap standard errors and $p$-values on the single BM task, with the nuisance imputation models fitted using KM. Through the output of DEFUSE, one can conclude that given other risk factors, a higher risk of HD is associated with low HDL at the level $0.1$ as well as type II diabetes in the level $0.05$. Notably, with $\mathcal{L}\mathcal{C}$-only and SSL, both the effects of HDL and diabetes are shown to be insignificant even in the $0.1$ level and diabetes is not significant in HTLGMM. This suggests that DEFUSE is more powerful than these methods to discover both signals. Interestingly, it is well-documented in biomedical literature on the associations between heart disease and diabetes (Haffner,, 2000; Peters et al.,, 2014, e.g.) as well as low HDL (Després et al.,, 2000; Rader and Hovingh,, 2014, e.g.), which supports the findings of DEFUSE.
At last, Table 9 summarizes the RE to the $\mathcal{L}\mathcal{C}$-only estimator in the multiple BM scenario introduced above. DEFUSE attains more than $25\%$ efficiency improvement over $\mathcal{L}\mathcal{C}$-only on all coefficients and more than $50\%$ variance reduction on that of age, which is again more significant than SSL. One can find the summary table of the estimates and $p$-values in this scenario in Appendix.
Method Age HDL Diabetes LRS
SemiEff (KM) 2.15 2.21 2.20 1.08
DEFUSE1 $\xi(\boldsymbol{x},y)=1$ (KM) 2.15 2.21 2.20 1.08
DEFUSE1 $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ (KM) 2.16 2.22 2.20 1.17
DEFUSE (KM) 2.19 2.24 2.22 1.24
SemiEff (Lasso) 2.14 2.20 2.19 0.97
DEFUSE1 $\xi(\boldsymbol{x},y)=1$ (Lasso) 2.14 2.20 2.19 1.00
DEFUSE1 $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ (Lasso) 2.14 2.20 2.19 1.10
DEFUSE (Lasso) 2.17 2.23 2.22 1.16
SSL 1.07 0.99 1.00 1.00
HTLGMM 1.76 1.55 1.56 1.02
Table 7: Relative efficiency (RE) to the standard $\mathcal{L}\mathcal{C}$-only estimator on each coefficient in the single BM scenario of our real example. KM stands for kernel machine and LRS for lab risk score.
@@
$\mathcal{L}\mathcal{C}$-only $0.41_{0.134}$ 0.002 $0.35_{0.232}$ 0.136 $0.15_{0.232}$ 0.509 $0.25_{0.116}$ 0.032
SemiEff $0.46_{0.092}$ $<$0.001 $0.26_{0.156}$ 0.093 $0.31_{0.156}$ 0.047 $0.32_{0.112}$ 0.004
DEFUSE1 $0.46_{0.091}$ $<$0.001 $0.27_{0.156}$ 0.088 $0.30_{0.156}$ 0.051 $0.32_{0.107}$ 0.002
DEFUSE $0.46_{0.091}$ $<$0.001 $0.28_{0.155}$ 0.07 $0.30_{0.155}$ 0.05 $0.39_{0.104}$ $<$0.001
SSL $0.41_{0.130}$ 0.001 $0.34_{0.233}$ 0.140 $0.16_{0.232}$ 0.480 $0.25_{0.116}$ 0.032
HTLGMM $0.47_{0.101}$ $<$0.001 $0.37_{0.186}$ 0.046 $0.20_{0.185}$ 0.174 $0.25_{0.115}$ 0.029
Table 8: Fitted coefficient (coe), empirical standard error (se), and $p$-value for the coefficient of each covariate in the single BM scenario of our real-world study. The